Just need help with these, thank you in advance!

1.Suppose that you invest $1,300 in stock. Four years later, your investment yields $1,515. What is the rate of return of your

2.You are going to buy a new car worth $25,700. The dealer computes your monthly payment to be $523.55 for 60 months of financi
Answer #1

(1)

Rate of return = ($1,515 / $1,300)(1/4) - 1 = (1.1654)0.25 - 1 = 1.039 - 1 = 0.039 = 3.9%

(2)

Monthly rate of return is found by Excel RATE function as follows.

Purchase Price/Loan Amount (PV) ($) = -25,700
Monthly Loan Payment (PMT) ($) = 523.55
Number of Months (NPER) = 60
Monthly Rate of Return = 0.68%

Effective rate = (1 + 0.0068)12 - 1 = (1.0068)12 - 1 = 1.085 - 1 = 0.085 = 8.5%